| Event Type | Date | Time | Source |
|---|---|---|---|
| Failure | 21-Jul-2001 | 17:00:00 | Logbook I, Page 7 |
| Run and bypass valve failure. Dislodged manually. | |||
| Insert Change | 21-Jul-2001 | 18:26:33 | HCLog 36553 |
| Insert ??? | Top: ??? (Deut) | Bot: ??? (Deut) | |
| Calibration | 22-Jul-2001 | 18:40:00 | Logbook I, Page 10 |
| LHe level vs. scattering rate study. | |||
| Calibration | 22-Jul-2001 | 23:38:32 | HCLog 36635 |
| Target position scan w/ beam. | |||
| Calibration | 23-Jul-2001 | 19:15:00 | Logbook I, Page 14 |
| LHe level vs. beam position study. | |||
| Calibration | 24-Jul-2001 | 04:14:00 | Logbook I, Page 18 |
| Target vs. Encoder position. Jon Farrell calculated. | |||
| Calibration | 23-Jul-2001 | 23:11:26 | Logbook I, Page 19 |
| Encoder position. Josh Pierce calculated. | |||
| Calibration | 24-Jul-2001 | 02:02:10 | HCLog 36693 |
| Densities for various materials. | |||
| Calibration | 24-Jul-2001 | 11:00:00 | Logbook I, Page 21 |
| Scattering rates vs. Encoder position. | |||
| Change | 24-Jul-2001 | 11:49:00 | Logbook I, Page 22 |
| New 3He manometer installed, old one removed. | |||
| Calibration | 24-Jul-2001 | 17:00:00 | Logbook I, Page 23 |
| Table to be positioned by encoder values. | |||
| Change | 24-Jul-2001 | 20:00:00 | Logbook I, Page 23 |
| Insert plunger pins changed from plastic to stainless. | |||
| Failure | 24-Jul-2001 | 22:30:00 | Logbook I, Page 23 |
| Bypass valve problem. Physically reset bypass valve. | |||
| Change | 24-Jul-2001 | 23:31:18 | HCLog 36747 |
| New encoder positions. | |||
| Calibration | 25-Jul-2001 | 15:00:00 | Logbook I, Page 28 |
| Insert pin position vs. insert angle plots. | |||
| Failure | 25-Jul-2001 | 18:32:00 | Logbook I, Page 29 |
| Magnet quench. | |||
| Failure | 26-Jul-2001 | 04:37:00 | Logbook I, Page 30 |
| Raster off w/ beam on bottom ND3 target. | |||
| Failure | 26-Jul-2001 | 22:41:48 | HCLog 36888 |
| Magnet quench. | |||
| Calibration | 26-Jul-2001 | 23:09:46 | HCLog 36889 |
| Encoder and target table values. | |||
| Calibration | 26-Jul-2001 | 23:50:32 | HCLog 36891 |
| Pictures of beam spot on ammonia. | |||
| Change | 27-Jul-2001 | 01:08:00 | Logbook I, Page 31 |
| Copper disk inserted in "empty" position. Also hclog 36890. | |||
| Calibration | 27-Jul-2001 | 05:00:00 | Logbook I, Page 32 |
| Chart for He3 vs. temperature. | |||
| Calibration | 28-Jul-2001 | 03:00:00 | Logbook I, Page 34 |
| Target position plot using copper insert. | |||
| Calibration | 29-Jul-2001 | 00:55:00 | Logbook I, Page 36 |
| Encoder values recalculated. | |||
| Change | 29-Jul-2001 | 03:45:00 | Logbook I, Page 37 |
| Windows based DAQ installed in SOS hut. | |||
| Change | 29-Jul-2001 | 04:30:26 | HCLog 37009 |
| New encoder values. | |||
| Change | 29-Jul-2001 | 18:22:00 | Logbook I, Page 38 |
| Cable changes to DIO box and DAQ computer. NMR retuned. | |||
| Change | 30-Jul-2001 | 08:00:00 | Logbook I, Page 40 |
| Cryostat moved by 4mm to recenter target with beam. | |||
| Change | 31-Jul-2001 | 08:44:00 | Logbook I, Page 41 |
| Terminators attached to unused SMA connectors on insert. | |||
| Change | 31-Jul-2001 | 11:00:00 | Logbook I, Page 42 |
| Replaced section of phase cable suspected to be bad. | |||
| Change | 31-Jul-2001 | 11:48:00 | Logbook I, Page 42 |
| Sampling frequency changed so noise and sweep are out of phase. | |||
| TE Calibration | 31-Jul-2001 | 14:04:10 | Logbook I, Page 42 |
| Deuteron | 20 samples | 1.553K | |
| TE Calibration | 31-Jul-2001 | 18:05:07 | Logbook I, Page 43 |
| Deuteron | 11 samples | 1.563K | |
| Calibration | 31-Jul-2001 | 23:00:00 | Logbook I, Page 46 |
| Peak asymmetry polarization calculated. | |||
| Calibration | 01-Aug-2001 | 11:50:00 | Logbook I, Page 47 |
| Gain setting were adjusted on Baselines. | |||
| Change | 01-Aug-2001 | 11:50:00 | Logbook I, Page 47 |
| RF power cable swapped on Q-meters for noise reasons. | |||
| Change | 01-Aug-2001 | 12:57:00 | Logbook I, Page 48 |
| Several cables changed on try to eliminate noise on bottom target. | |||
| Calibration | 01-Aug-2001 | 13:55:00 | Logbook I, Page 48 |
| Switches for target position adjusted. | |||
| Failure | 02-Aug-2001 | 01:51:00 | Logbook I, Page 53 |
| Pumps tripped off. | |||
| Failure | 02-Aug-2001 | 09:10:00 | Logbook I, Page 55 |
| Pumps tripped off. | |||
| Calibration | 02-Aug-2001 | 16:04:47 | HCLog 37254 |
| Materials composing target. | |||
| Anneal | 02-Aug-2001 | 20:34:00 | Logbook I, Page 56 |
| 80K | 45 minutes | ||
| Anneal | 04-Aug-2001 | 09:59:48 | Logbook I, Page 61 |
| 80K | 45 minutes | ||
| Change | 04-Aug-2001 | 12:00:00 | Logbook I, Page 61 |
| Qmeters retuned. | |||
| TE Calibration | 04-Aug-2001 | 14:31:00 | Logbook I, Page 62 |
| Deuteron | 15 samples | 1.574K | |
| Change | 04-Aug-2001 | 17:00:00 | Logbook I, Page 63 |
| Phase cable switched on bottom cell. | |||
| TE Calibration | 04-Aug-2001 | 17:33:27 | Logbook I, Page 63 |
| Deuteron | 16 samples | 1.588K | |
| TE Calibration | 05-Aug-2001 | 20:59:00 | Logbook I, Page 67 |
| Deuteron | 15 samples | 1.677K | |
| Failure | 06-Aug-2001 | 09:50:00 | Logbook I, Page 70 |
| Separator valve not working. Computer rebooted. | |||
| Anneal | 06-Aug-2001 | 11:01:00 | Logbook I, Page 70 |
| Empty Target | 80K | 35 minutes | |
| Change | 07-Aug-2001 | 09:55:00 | Logbook I, Page 76 |
| Labview restarted to correct spurious polarization reading. | |||
| Anneal | 08-Aug-2001 | 09:45:00 | Logbook I, Page 82 |
| Empty Target | 60K | 45 minutes | |
| Anneal | 08-Aug-2001 | 18:00:00 | Logbook I, Page 84 |
| Empty Target | 80K | 10 minutes | |
| Anneal | 09-Aug-2001 | 18:20:00 | Logbook I, Page 92 |
| Empty Target | 80K | 25 minutes | |
| Anneal | 10-Aug-2001 | 17:28:00 | Logbook I, Page 100 |
| Empty Target | 90K | 45 minutes | |
| Anneal | 12-Aug-2001 | 08:21:00 | Logbook I, Page 106 |
| 85K | 39 minutes | ||
| Calibration | 13-Aug-2001 | 04:30:00 | Logbook I, Page 110 |
| Polarization decay study. | |||
| Failure | 14-Aug-2001 | 09:45:00 | Logbook I, Page 112 |
| Safety crew forced restricted access - ramp down magnet. | |||
| Insert Change | 15-Aug-2001 | 10:00:00 | Logbook I, Page 113 |
| Insert C | Top: Gen01_Bottle_3 (Deut) | Bot: 95%_Gen01_Bottle_1,5%_Gen01_Bottle_2 (Deut) | |
| TE Calibration | 16-Aug-2001 | 00:16:00 | Logbook I, Page 114 |
| Deuteron | 21 samples | 1.617K | |
| TE Calibration | 16-Aug-2001 | 01:48:00 | Logbook I, Page 115 |
| Deuteron | 22 samples | 1.629K | |
| Failure | 16-Aug-2001 | 05:00:00 | Logbook I, Page 116 |
| Pump trip. | |||
| Failure | 17-Aug-2001 | 13:15:00 | Logbook I, Page 122 |
| Separator valve stuck. | |||
| Anneal | 18-Aug-2001 | 12:00:00 | Logbook I, Page 127 |
| 0K | 0 minutes | ||
| Calibration | 19-Aug-2001 | 16:30:00 | Logbook I, Page 133 |
| Lambda/2 cable lengths: top=184cm, bot=191cm, top phase=76cm, v=0.78 | |||
| Anneal | 19-Aug-2001 | 16:38:00 | Logbook I, Page 133 |
| 82K | 35 minutes | ||
| Failure | 19-Aug-2001 | 18:24:00 | Logbook I, Page 133 |
| Pump trip. | |||
| TE Calibration | 21-Aug-2001 | 10:48:00 | Logbook I, Page 144 |
| Deuteron | 19 samples | 1.635K | |
| TE Calibration | 21-Aug-2001 | 13:23:00 | Logbook I, Page 144 |
| Deuteron | 20 samples | 1.639K | |
| TE Calibration | 21-Aug-2001 | 15:43:00 | Logbook I, Page 144 |
| Deuteron | 17 samples | 1.638K | |
| TE Calibration | 21-Aug-2001 | 16:35:00 | Logbook I, Page 144 |
| Deuteron | 20 samples | 1.647K | |
| Anneal | 21-Aug-2001 | 18:19:00 | Logbook I, Page 147 |
| Empty Target | 95K | 15 minutes | |
| Anneal | 22-Aug-2001 | 23:59:00 | Logbook I, Page 152 |
| Empty Target | 95K | 30 minutes | |
| Anneal | 24-Aug-2001 | 10:21:00 | Logbook I, Page 157 |
| 95K | 45 minutes | ||
| Failure | 27-Aug-2001 | 00:56:00 | Logbook II, Page 13 |
| Pump fuse blown while preparing for anneal. | |||
| Anneal | 27-Aug-2001 | 01:06:00 | Logbook II, Page 13 |
| 95K | 60 minutes | ||
| Anneal | 28-Aug-2001 | 21:30:00 | Logbook II, Page 20 |
| 95K | 60 minutes | ||
| Failure | 29-Aug-2001 | 06:25:00 | Logbook II, Page 22 |
| Pumps tripped off | |||
| Calibration | 31-Aug-2001 | 07:03:00 | Logbook II, Page 32 |
| Possible Decay measurement. | |||
| Change | 31-Aug-2001 | 07:35:01 | Logbook II, Page 32 |
| He3 Manometer removed. | |||
| TE Calibration | 31-Aug-2001 | 12:31:02 | Logbook II, Page 34 |
| Deuteron | 20 samples | 1.413K | |
| TE Calibration | 31-Aug-2001 | 16:08:41 | Logbook II, Page 35 |
| Deuteron | 20 samples | 1.418K | |
| TE Calibration | 31-Aug-2001 | 22:21:44 | Logbook II, Page 37 |
| Deuteron | 20 samples | 1.423K | |
| TE Calibration | 01-Sep-2001 | 03:24:00 | Logbook II, Page 38 |
| Deuteron | 19 samples | 1.592K | |
| TE Calibration | 01-Sep-2001 | 06:08:00 | Logbook II, Page 39 |
| Deuteron | 19 samples | 1.639K | |
| TE Calibration | 01-Sep-2001 | 10:09:22 | Logbook II, Page 40 |
| Deuteron | 20 samples | 1.648K | |
| TE Calibration | 01-Sep-2001 | 12:41:55 | Logbook II, Page 41 |
| Deuteron | 19 samples | 1.658K | |
| TE Calibration | 01-Sep-2001 | 14:03:30 | Logbook II, Page 41 |
| Deuteron | 20 samples | 1.653K | |
| TE Calibration | 01-Sep-2001 | 21:56:00 | Logbook II, Page 44 |
| Deuteron | 10 samples | 1.643K | |
| TE Calibration | 01-Sep-2001 | 23:04:00 | Logbook II, Page 44 |
| Deuteron | 15 samples | 1.637K | |
| TE Calibration | 02-Sep-2001 | 00:59:00 | Logbook II, Page 45 |
| Deuteron | 10 samples | 1.641K | |
| TE Calibration | 02-Sep-2001 | 03:15:00 | Logbook II, Page 46 |
| Deuteron | 13 samples | 1.641K | |
| Change | 02-Sep-2001 | 04:28:00 | Logbook II, Page 46 |
| Scale on He4 manometer changed to accommodate 1.9 K TE. | |||
| TE Calibration | 02-Sep-2001 | 05:08:00 | Logbook II, Page 47 |
| Deuteron | 12 samples | 1.876K | |
| TE Calibration | 02-Sep-2001 | 07:43:51 | Logbook II, Page 48 |
| Deuteron | 20 samples | 1.878K | |
| TE Calibration | 02-Sep-2001 | 09:18:46 | Logbook II, Page 49 |
| Deuteron | 20 samples | 1.881K | |
| TE Calibration | 02-Sep-2001 | 12:04:00 | Logbook II, Page 50 |
| Deuteron | 16 samples | 1.880K | |
| TE Calibration | 02-Sep-2001 | 13:21:00 | Logbook II, Page 50 |
| Deuteron | 10 samples | 1.887K | |
| TE Calibration | 02-Sep-2001 | 14:17:00 | Logbook II, Page 51 |
| Deuteron | 20 samples | 1.895K | |
| Anneal | 04-Sep-2001 | 16:26:00 | Logbook II, Page 52 |
| Empty Target | 95K | 60 minutes | |
| Anneal | 07-Sep-2001 | 15:25:00 | Logbook II, Page 67 |
| Empty Target | 95K | 45 minutes | |
| Anneal | 08-Sep-2001 | 23:40:00 | Logbook II, Page 72 |
| Empty Target | 95K | 60 minutes | |
| Failure | 09-Sep-2001 | 05:24:00 | Logbook II, Page 73 |
| Run valve stuck. | |||
| Change | 09-Sep-2001 | 19:00:00 | HCLog 40359 |
| EPICS data added to the event stream via EDR.vi | |||
| Anneal | 10-Sep-2001 | 09:18:00 | Logbook II, Page 78 |
| 95K | 60 minutes | ||
| Anneal | 11-Sep-2001 | 11:48:00 | Logbook II, Page 83 |
| Empty Target | 95K | 60 minutes | |
| Failure | 13-Sep-2001 | 17:10:00 | Logbook II, Page 88 |
| Target mover problem. | |||
| Anneal | 13-Sep-2001 | 17:25:00 | Logbook II, Page 89 |
| Empty Target | 95K | 90 minutes | |
| Failure | 14-Sep-2001 | 08:00:00 | Logbook II, Page 90 |
| Target mover problem. | |||
| Change | 15-Sep-2001 | 00:03:00 | Logbook II, Page 92 |
| One month shut down preparations. | |||
| TE Calibration | 17-Sep-2001 | 02:45:11 | Logbook II, Page 94 |
| Deuteron | 21 samples | 1.532K | |
| TE Calibration | 17-Sep-2001 | 05:16:20 | Logbook II, Page 95 |
| Deuteron | 22 samples | 1.537K | |
| TE Calibration | 17-Sep-2001 | 08:12:13 | Logbook II, Page 96 |
| Deuteron | 10 samples | 1.531K | |
| TE Calibration | 17-Sep-2001 | 09:14:22 | Logbook II, Page 96 |
| Deuteron | 10 samples | 1.546K | |
| TE Calibration | 17-Sep-2001 | 10:35:05 | Logbook II, Page 97 |
| Deuteron | 11 samples | 1.549K | |
| TE Calibration | 17-Sep-2001 | 12:03:36 | Logbook II, Page 97 |
| Deuteron | 10 samples | 1.550K | |
| TE Calibration | 17-Sep-2001 | 13:36:24 | Logbook II, Page 98 |
| Deuteron | 10 samples | 1.554K | |
| TE Calibration | 17-Sep-2001 | 16:00:14 | Logbook II, Page 98 |
| Deuteron | 11 samples | 1.560K | |
| TE Calibration | 17-Sep-2001 | 17:27:58 | Logbook II, Page 99 |
| Deuteron | 10 samples | 1.543K | |
| Anneal | 17-Sep-2001 | 18:33:00 | Logbook II, Page 99 |
| 95K | 60 minutes | ||
| Calibration | 18-Sep-2001 | 07:52:00 | Logbook II, Page 103 |
| Polarization decay study. | |||
| Insert Change | 18-Sep-2001 | 14:00:00 | Logbook II, Page 104 |
| Insert C | Top: xxx (Deut) | Bot: xxx (Deut) | |
| Calibration | 18-Sep-2001 | 15:00:00 | Logbook II, Page 104 |
| Compared He3 to He4 manometer. | |||
| Change | 19-Sep-2001 | 16:00:00 | Logbook II, Page 107 |
| Shut down. | |||
| Insert Change | 14-Oct-2001 | 15:00:00 | Logbook II, Page 109 |
| Insert x | Top: xxx (Deut) | Bot: xxx (Deut) | |
| Calibration | 14-Oct-2001 | 15:01:00 | Logbook II, Page 109 |
| He3 to He4 calibration. | |||
| Change | 15-Oct-2001 | 12:00:00 | Logbook II, Page 110 |
| He3 and He4 manometers adjusted. | |||
| Insert Change | 15-Oct-2001 | 20:00:00 | Logbook II, Page 112 |
| Insert B | Top: E155_Bottle_4_+_Gen_Bottle_1 (Deut) | Bot: 6/7/01_Consol_+_Gen_Bottle_1 (Deut) | |
| TE Calibration | 17-Oct-2001 | 11:12:37 | Logbook II, Page 113 |
| Deuteron | 20 samples | 1.543K | |
| TE Calibration | 17-Oct-2001 | 14:34:12 | Logbook II, Page 114 |
| Deuteron | 20 samples | 1.528K | |
| TE Calibration | 22-Oct-2001 | 16:28:55 | Logbook II, Page 114 |
| Deuteron | 20 samples | 1.528K | |
| Calibration | 22-Oct-2001 | 12:00:00 | Logbook II, Page 116 |
| He4 and He3 temperature chart. | |||
| Change | 22-Oct-2001 | 15:07:00 | Logbook II, Page 117 |
| Wiring chart for MIO card connections. | |||
| TE Calibration | 22-Oct-2001 | 18:08:08 | Logbook II, Page 118 |
| Deuteron | 10 samples | 1.584K | |
| TE Calibration | 22-Oct-2001 | 19:34:26 | Logbook II, Page 119 |
| Deuteron | 10 samples | 1.577K | |
| TE Calibration | 22-Oct-2001 | 20:58:22 | Logbook II, Page 120 |
| Deuteron | 10 samples | 1.573K | |
| TE Calibration | 22-Oct-2001 | 22:20:29 | Logbook II, Page 120 |
| Deuteron | 10 samples | 1.668K | |
| TE Calibration | 22-Oct-2001 | 23:45:47 | Logbook II, Page 122 |
| Deuteron | 5 samples | 1.563K | |
| TE Calibration | 23-Oct-2001 | 00:36:11 | Logbook II, Page 123 |
| Deuteron | 10 samples | 1.552K | |
| TE Calibration | 23-Oct-2001 | 02:07:51 | Logbook II, Page 124 |
| Deuteron | 10 samples | 1.555K | |
| TE Calibration | 23-Oct-2001 | 03:34:40 | Logbook II, Page 125 |
| Deuteron | 10 samples | 1.558K | |
| TE Calibration | 23-Oct-2001 | 05:03:11 | Logbook II, Page 126 |
| Deuteron | 10 samples | 1.570K | |
| TE Calibration | 23-Oct-2001 | 06:24:05 | Logbook II, Page 127 |
| Deuteron | 11 samples | 1.575K | |
| TE Calibration | 23-Oct-2001 | 07:47:23 | Logbook II, Page 128 |
| Deuteron | 11 samples | 1.575K | |
| TE Calibration | 23-Oct-2001 | 11:10:49 | Logbook II, Page 130 |
| Deuteron | 10 samples | 1.573K | |
| TE Calibration | 23-Oct-2001 | 12:29:39 | Logbook II, Page 131 |
| Deuteron | 10 samples | 1.578K | |
| TE Calibration | 23-Oct-2001 | 13:56:28 | Logbook II, Page 132 |
| Deuteron | 13 samples | 1.555K | |
| TE Calibration | 23-Oct-2001 | 15:54:55 | Logbook II, Page 133 |
| Deuteron | 10 samples | 1.562K | |
| TE Calibration | 23-Oct-2001 | 17:11:44 | Logbook II, Page 134 |
| Deuteron | 11 samples | 1.566K | |
| TE Calibration | 23-Oct-2001 | 18:38:16 | Logbook II, Page 134 |
| Deuteron | 10 samples | 1.565K | |
| TE Calibration | 23-Oct-2001 | 20:11:25 | Logbook II, Page 136 |
| Deuteron | 11 samples | 1.564K | |
| TE Calibration | 23-Oct-2001 | 21:41:42 | Logbook II, Page 137 |
| Deuteron | 10 samples | 1.561K | |
| TE Calibration | 23-Oct-2001 | 23:07:38 | Logbook II, Page 138 |
| Deuteron | 8 samples | 1.561K | |
| TE Calibration | 24-Oct-2001 | 04:48:48 | Logbook II, Page 140 |
| Deuteron | 10 samples | 1.303K | |
| TE Calibration | 24-Oct-2001 | 06:17:41 | Logbook II, Page 141 |
| Deuteron | 10 samples | 1.287K | |
| TE Calibration | 24-Oct-2001 | 07:39:35 | Logbook II, Page 144 |
| Deuteron | 11 samples | 1.287K | |
| TE Calibration | 24-Oct-2001 | 09:07:36 | Logbook II, Page 146 |
| Deuteron | 12 samples | 1.286K | |
| TE Calibration | 24-Oct-2001 | 14:16:05 | Logbook II, Page 147 |
| Deuteron | 10 samples | 1.630K | |
| Calibration | 24-Oct-2001 | 17:45:00 | Logbook II, Page 148 |
| Top target baseline with field off. | |||
| Calibration | 24-Oct-2001 | 19:55:00 | Logbook II, Page 148 |
| Manometer field sensitivity measurement. | |||
| Failure | 27-Oct-2001 | 16:30:00 | Logbook II, Page 152 |
| Fuse in microwave switch blown. | |||
| Failure | 28-Oct-2001 | 02:19:00 | Logbook II, Page 153 |
| Pump trip | |||
| Failure | 28-Oct-2001 | 09:30:00 | Logbook II, Page 154 |
| Temperature readings failed. | |||
| Failure | 28-Oct-2001 | 15:30:00 | Logbook II, Page 155 |
| Microwave switch malfunctioning. | |||
| Anneal | 29-Oct-2001 | 13:25:00 | Logbook II, Page 160 |
| 85K | 30 minutes | ||
| TE Calibration | 30-Oct-2001 | 14:39:46 | Logbook II, Page 163 |
| Deuteron | 10 samples | 1.534K | |
| Failure | 31-Oct-2001 | 19:38:00 | Logbook II, Page 166 |
| Microwave switch not working. | |||
| Anneal | 01-Nov-2001 | 10:06:00 | Logbook II, Page 168 |
| Carbon Target | 80K | 60 minutes | |
| Change | 02-Nov-2001 | 10:00:00 | Logbook II, Page 173 |
| Bottom deuteron taken at gain 50. | |||
| Change | 02-Nov-2001 | 10:15:00 | Logbook II, Page 173 |
| Microwave switch removed and replaced with strait waveguide. | |||
| Anneal | 02-Nov-2001 | 11:35:00 | Logbook II, Page 174 |
| 80K | 10 minutes | ||
| Anneal | 03-Nov-2001 | 16:09:00 | Logbook II, Page 179 |
| Empty Target | 65K | 15 minutes | |
| Anneal | 05-Nov-2001 | 01:01:00 | Logbook II, Page 185 |
| Empty Target | 65K | 5 minutes | |
| Anneal | 05-Nov-2001 | 17:11:00 | Logbook II, Page 189 |
| Carbon Target | 75K | 10 minutes | |
| Failure | 05-Nov-2001 | 17:25:00 | Logbook II, Page 190 |
| Run valve stuck due to ice. | |||
| Change | 06-Nov-2001 | 10:00:00 | Logbook II, Page 192 |
| MIO card in computer replaced. | |||
| TE Calibration | 06-Nov-2001 | 12:51:47 | Logbook II, Page 193 |
| Deuteron | 12 samples | 1.499K | |
| TE Calibration | 06-Nov-2001 | 14:29:31 | Logbook II, Page 194 |
| Deuteron | 10 samples | 1.474K | |
| Anneal | 06-Nov-2001 | 16:56:00 | Logbook II, Page 195 |
| Empty Target | 80K | 10 minutes | |
| Anneal | 07-Nov-2001 | 10:08:00 | Logbook II, Page 200 |
| Empty Target | 85K | 10 minutes | |
| Anneal | 08-Nov-2001 | 11:01:00 | Logbook II, Page 207 |
| Empty Target | 95K | 15 minutes | |
| Anneal | 09-Nov-2001 | 22:52:00 | Logbook II, Page 213 |
| Carbon Target | 95K | 30 minutes | |
| Failure | 10-Nov-2001 | 02:30:00 | Logbook II, Page 214 |
| Run valve stuck. | |||
| Anneal | 11-Nov-2001 | 00:37:00 | Logbook II, Page 217 |
| Empty Target | 95K | 45 minutes | |
| Anneal | 12-Nov-2001 | 08:54:00 | Logbook II, Page 221 |
| 95K | 60 minutes | ||
| Anneal | 13-Nov-2001 | 17:57:00 | Logbook II, Page 226 |
| Carbon Target | 95K | 60 minutes | |
| Anneal | 14-Nov-2001 | 17:09:00 | Logbook II, Page 230 |
| Empty Target | 95K | 51 minutes | |
| Anneal | 15-Nov-2001 | 19:37:00 | Logbook II, Page 235 |
| Empty Target | 95K | 60 minutes | |
| Anneal | 16-Nov-2001 | 21:08:00 | Logbook II, Page 240 |
| Empty Target | 95K | 60 minutes | |
| Failure | 18-Nov-2001 | 01:25:00 | Logbook II, Page 245 |
| Computer crash during anneal. | |||
| Anneal | 18-Nov-2001 | 01:30:00 | Logbook II, Page 246 |
| 95K | 30 minutes | ||
| Anneal | 18-Nov-2001 | 23:45:00 | Logbook II, Page 256 |
| Empty Target | 95K | 75 minutes | |
| Anneal | 19-Nov-2001 | 23:30:00 | Logbook II, Page 260 |
| Empty Target | 95K | 60 minutes | |
| Failure | 20-Nov-2001 | 17:20:00 | Logbook II, Page 264 |
| Separator valve stuck. | |||
| Anneal | 20-Nov-2001 | 18:38:00 | Logbook II, Page 264 |
| Empty Target | 95K | 60 minutes | |
| Change | 21-Nov-2001 | 08:00:00 | Logbook II, Page 269 |
| Preparations for holiday shutdown made. | |||
| TE Calibration | 26-Nov-2001 | 14:12:43 | Logbook II, Page 270 |
| Deuteron | 10 samples | 1.538K | |
| TE Calibration | 26-Nov-2001 | 15:42:02 | Logbook II, Page 271 |
| Deuteron | 10 samples | 1.554K | |
| TE Calibration | 26-Nov-2001 | 17:11:17 | Logbook II, Page 272 |
| Deuteron | 10 samples | 1.536K | |
| TE Calibration | 26-Nov-2001 | 18:51:50 | Logbook II, Page 272 |
| Deuteron | 10 samples | 1.550K | |
| TE Calibration | 26-Nov-2001 | 20:07:05 | Logbook II, Page 273 |
| Deuteron | 10 samples | 1.551K | |
| TE Calibration | 26-Nov-2001 | 21:19:52 | Logbook II, Page 274 |
| Deuteron | 10 samples | 1.547K | |
| TE Calibration | 26-Nov-2001 | 22:38:41 | Logbook II, Page 274 |
| Deuteron | 10 samples | 1.542K | |
| TE Calibration | 26-Nov-2001 | 23:56:35 | Logbook II, Page 275 |
| Deuteron | 10 samples | 1.546K | |
| TE Calibration | 27-Nov-2001 | 01:18:18 | Logbook II, Page 275 |
| Deuteron | 10 samples | 1.551K | |
| TE Calibration | 27-Nov-2001 | 02:44:49 | Logbook II, Page 276 |
| Deuteron | 10 samples | 1.573K | |
| TE Calibration | 27-Nov-2001 | 04:07:19 | Logbook II, Page 276 |
| Deuteron | 10 samples | 1.546K | |
| TE Calibration | 27-Nov-2001 | 05:29:08 | Logbook II, Page 277 |
| Deuteron | 10 samples | 1.554K | |
| TE Calibration | 27-Nov-2001 | 06:52:35 | Logbook II, Page 277 |
| Deuteron | 10 samples | 1.554K | |
| Change | 27-Nov-2001 | 08:10:00 | Logbook II, Page 278 |
| Gain 50 used for baselines in following TEs. | |||
| TE Calibration | 27-Nov-2001 | 08:25:32 | Logbook II, Page 278 |
| Deuteron | 10 samples | 1.553K | |
| TE Calibration | 27-Nov-2001 | 10:02:11 | Logbook II, Page 279 |
| Deuteron | 10 samples | 1.538K | |
| TE Calibration | 27-Nov-2001 | 11:30:22 | Logbook II, Page 279 |
| Deuteron | 10 samples | 1.548K | |
| Failure | 27-Nov-2001 | 13:35:00 | Logbook II, Page 280 |
| Material fell out of cups. | |||
| Insert Change | 27-Nov-2001 | 15:00:00 | Logbook II, Page 281 |
| Insert C | Top: 50%_Gen01_Bottle_1,50%_Gen01_Bottle_2 (Deut) | Bot: 66%_Gen01_Bottle_5,14%_Gen01_Bottle_6,20%_Gen01_Bottle_2 (Deut) | |
| TE Calibration | 27-Nov-2001 | 20:19:12 | Logbook II, Page 282 |
| Deuteron | 10 samples | 1.540K | |
| TE Calibration | 27-Nov-2001 | 21:34:37 | Logbook II, Page 283 |
| Deuteron | 10 samples | 1.539K | |
| TE Calibration | 27-Nov-2001 | 22:48:20 | Logbook II, Page 283 |
| Deuteron | 10 samples | 1.543K | |
| TE Calibration | 28-Nov-2001 | 00:50:47 | Logbook II, Page 284 |
| Deuteron | 10 samples | 1.559K | |
| TE Calibration | 28-Nov-2001 | 02:09:16 | Logbook II, Page 284 |
| Deuteron | 10 samples | 1.534K | |
| TE Calibration | 28-Nov-2001 | 03:32:26 | Logbook II, Page 285 |
| Deuteron | 10 samples | 1.539K | |
| TE Calibration | 28-Nov-2001 | 05:38:01 | Logbook II, Page 285 |
| Deuteron | 10 samples | 1.554K | |
| TE Calibration | 28-Nov-2001 | 07:16:24 | Logbook II, Page 286 |
| Deuteron | 10 samples | 1.557K | |
| TE Calibration | 28-Nov-2001 | 08:54:02 | Logbook II, Page 286 |
| Deuteron | 10 samples | 1.517K | |
| TE Calibration | 28-Nov-2001 | 11:34:48 | Logbook II, Page 287 |
| Deuteron | 10 samples | 1.527K | |
| Failure | 28-Nov-2001 | 18:15:00 | Logbook II, Page 287 |
| Microwave switch malfunctioning. | |||
| Failure | 28-Nov-2001 | 21:30:00 | Logbook II, Page 288 |
| Magnet quench. | |||
| Calibration | 29-Nov-2001 | 01:23:00 | Logbook II, Page 289 |
| Relaxation study. | |||
| Calibration | 30-Nov-2001 | 15:00:00 | Logbook II, Page 289 |
| Beam found to be uncentered. | |||
| Change | 01-Dec-2001 | 00:05:00 | Logbook II, Page 289 |
| Encoder values changed. | |||
| Anneal | 01-Dec-2001 | 21:29:00 | Logbook II, Page 295 |
| 95K | 60 minutes | ||
| Anneal | 02-Dec-2001 | 21:51:00 | Logbook II, Page 299 |
| Empty Target | 100K | 60 minutes | |
| Failure | 04-Dec-2001 | 08:25:00 | Logbook III, Page 1 |
| GPIB interface changed out. | |||
| TE Calibration | 04-Dec-2001 | 10:55:16 | Logbook III, Page 2 |
| Deuteron | 9 samples | 1.510K | |
| TE Calibration | 04-Dec-2001 | 12:25:13 | Logbook III, Page 3 |
| Deuteron | 20 samples | 1.542K | |
| TE Calibration | 04-Dec-2001 | 13:48:22 | Logbook III, Page 3 |
| Deuteron | 10 samples | 1.511K | |
| TE Calibration | 04-Dec-2001 | 15:25:31 | Logbook III, Page 4 |
| Deuteron | 10 samples | 1.532K | |
| TE Calibration | 04-Dec-2001 | 17:45:49 | Logbook III, Page 5 |
| Deuteron | 10 samples | 1.518K | |
| TE Calibration | 04-Dec-2001 | 19:02:54 | Logbook III, Page 5 |
| Deuteron | 10 samples | 1.522K | |
| Anneal | 04-Dec-2001 | 20:50:00 | Logbook III, Page 6 |
| 95K | 60 minutes | ||
| Anneal | 08-Dec-2001 | 01:43:00 | Logbook III, Page 15 |
| Empty Target | 95K | 60 minutes | |
| Failure | 08-Dec-2001 | 03:20:00 | Logbook III, Page 15 |
| Run valve stuck. | |||
| Failure | 08-Dec-2001 | 17:05:00 | Logbook III, Page 17 |
| run valve stuck. | |||
| Anneal | 09-Dec-2001 | 05:16:00 | Logbook III, Page 19 |
| Carbon Target | 100K | 60 minutes | |
| Anneal | 10-Dec-2001 | 06:54:00 | Logbook III, Page 24 |
| 100K | 60 minutes | ||
| Failure | 10-Dec-2001 | 11:30:00 | Logbook III, Page 25 |
| Power tripped to target racks. | |||
| Failure | 11-Dec-2001 | 08:02:00 | Logbook III, Page 29 |
| Power tripped to target racks. | |||
| Calibration | 11-Dec-2001 | 08:15:00 | Logbook III, Page 29 |
| Relaxation study. | |||
| TE Calibration | 11-Dec-2001 | 15:40:05 | Logbook III, Page 30 |
| Deuteron | 9 samples | 1.531K | |
| TE Calibration | 11-Dec-2001 | 17:00:31 | Logbook III, Page 31 |
| Deuteron | 11 samples | 1.522K | |
| TE Calibration | 11-Dec-2001 | 20:09:14 | Logbook III, Page 32 |
| Deuteron | 5 samples | 1.540K | |
| TE Calibration | 11-Dec-2001 | 21:25:48 | Logbook III, Page 32 |
| Deuteron | 10 samples | 1.554K | |
| TE Calibration | 11-Dec-2001 | 22:49:45 | Logbook III, Page 34 |
| Deuteron | 10 samples | 1.551K | |
| Anneal | 12-Dec-2001 | 01:04:00 | Logbook III, Page 56 |
| Empty Target | 100K | 60 minutes | |
| Anneal | 13-Dec-2001 | 01:50:00 | Logbook III, Page 42 |
| Empty Target | 100K | 60 minutes | |
| Anneal | 14-Dec-2001 | 07:49:00 | Logbook III, Page 49 |
| 105K | 60 minutes | ||
| Anneal | 15-Dec-2001 | 10:43:00 | Logbook III, Page 55 |
| Top Target | 105K | 60 minutes | |
| Anneal | 16-Dec-2001 | 01:40:00 | Logbook III, Page 57 |
| Empty Target | 105K | 30 minutes | |
| Failure | 16-Dec-2001 | 06:48:00 | Logbook III, Page 57 |
| Run valve stuck. | |||
| Anneal | 16-Dec-2001 | 22:34:00 | Logbook III, Page 61 |
| 105K | 60 minutes | ||
| Anneal | 17-Dec-2001 | 17:00:00 | Logbook III, Page 64 |
| Top Target | 105K | 60 minutes | |
| Calibration | 18-Dec-2001 | 04:00:00 | Logbook III, Page 66 |
| Decay study. | |||
| TE Calibration | 18-Dec-2001 | 14:50:33 | Logbook III, Page 68 |
| Deuteron | 7 samples | 1.547K | |
| TE Calibration | 18-Dec-2001 | 16:06:15 | Logbook III, Page 69 |
| Deuteron | 11 samples | 1.569K | |
| TE Calibration | 18-Dec-2001 | 17:27:50 | Logbook III, Page 69 |
| Deuteron | 10 samples | 1.567K | |
| TE Calibration | 18-Dec-2001 | 20:35:19 | Logbook III, Page 71 |
| Deuteron | 10 samples | 1.517K | |
| TE Calibration | 18-Dec-2001 | 22:01:49 | Logbook III, Page 72 |
| Deuteron | 10 samples | 1.552K | |
| Anneal | 19-Dec-2001 | 00:15:00 | Logbook III, Page 73 |
| Top Target | 110K | 60 minutes | |
| Anneal | 19-Dec-2001 | 23:30:00 | Logbook III, Page 78 |
| Empty Target | 110K | 60 minutes | |
| Anneal | 20-Dec-2001 | 17:33:00 | Logbook III, Page 82 |
| Empty Target | 115K | 60 minutes | |
| Calibration | 21-Dec-2001 | 00:46:00 | Logbook III, Page 84 |
| Polarization dependence on scan steps and scan frequency. | |||
| Change | 21-Dec-2001 | 06:00:00 | Logbook III, Page 85 |
| Shutdown. | |||